OpenStudy (anonymous):

If (x + 5) - 2(4x - 1) = 0, what is the value of x? i got a decimal. how do i solve this?

OpenStudy (ivettef365):

you have x + 5 -8x + 2 = 0

OpenStudy (anonymous):

(x + 5) - 2(4x - 1) = 0 Distribute the -2 to the right side. (x + 5) - 8x + 1 = 0 Remove() x + 5 - 8x + 1 = 0 Add and Subtract.

OpenStudy (anonymous):

how do i remove the 0? @some_someone

OpenStudy (anonymous):

(x + 5) - 2(4x - 1) = 0 Distribute the -2 to the right side. (x + 5) - 8x + 1 = 0 Remove() x + 5 - 8x + 1 = 0 Add and Subtract. x - 8x = ___ 5 + 1 = ___

OpenStudy (anonymous):

I never said 0 @GABY11 I said remove() As in () <=== parenthesis

OpenStudy (anonymous):

wait how did you get positive 1? isnt it suppose to be positive 2? @some_someone

OpenStudy (ivettef365):

yes it is +2

OpenStudy (ivettef365):

-7x +7 = 0 -7x = -7 x = 1